(Please show calculations for study purposes)

Logistics Solutions provides order fulfillment services for dot.com merchants. The company maintains warehouses that stock items carried by its dot.com clients. When a client receives an order from a customer, the order is forwarded to Logistics Solutions, which pulls the item from storage, packs it, and ships it to the customer. The company uses a predetermined variable overhead rate based on direct labor-hours.

In the most recent month, 205,000 items were shipped to customers using 9,200 direct labor-hours. The company incurred a total of $33,580 in variable overhead costs.

According to the company’s standards, 0.02 direct labor-hours are required to fulfill an order for one item and the variable overhead rate is $3.70 per direct labor-hour.

Required:

1. What is the standard labor-hours allowed (SH) to ship 205,000 items to customers?

2. What is the standard variable overhead cost allowed (SH × SR) to ship 205,000 items to customers?

3. What is the variable overhead spending variance?

4. What is the variable overhead rate variance and the variable overhead efficiency variance?
Answer #1

Standard hours allowed

= 205,000*0.02

= 4100

Standard cost allowed = 4100*3.70

= 15,170

Voh spending variance = 15,170 - 33,580

= 18,410 U

Voh rate variance = (SR-AR) *AH

= (3.70 - 33580/9200)*9200

= 460 F

Voh efficiency variance = (SH-AH) *SR

= (4100-9200)*3.70

= 18,870 U